OpenStudy (anonymous):

Two cars travel at 40 and 60 mph, respectively. If the second car starts out 5 miles behind the first car, how many hours will it take the second car to overtake the first car?

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ \normalsize \text{let the hrs taken be}~~ x ~~\text{hrs , so } \hspace{.33em}\\~\\ 40x+5=60x }\end{align}\) solve for \(\large \begin{align} \color{black}{ x \hspace{.33em}\\~\\ }\end{align}\)